# Hierarchical Prompting Taxonomy: A Universal Evaluation Framework for Large Language Models

### v2

@@ 3.4.2 HP-Score for Manual HPF
where $\text{hp}_{\text{j}}$ represents the HP-Score of sample $j$, x represents the level of HPF at which task is addressed, $m$ represents the total number of levels in HPF, and $n$ denotes the total number of samples in the evaluation set.
3.4.3 HP-Score for Adaptive HPF
The prompt-selector can dynamically select the most suitable prompting strategy for a given task’s complexity from the manual HPF’s collection of prompting strategies. In order to determine the most effective prompting strategy to address the task, the prompt-selector was given a maximum number of iterations equivalent to the number of levels in the manual HPF. The score for ith iteration is i + x, where x is the level of the prompting strategy selected by the prompt-selector. If the LLM fails to address the task after all iterations, it is assigned a penalty corresponding to it’s dataset HP-Score. If LLM was able to address the task at ith iteration:
$\text{hp}_{\text{j}}=x+i$ (6)
If LLM was unable to address the task after all m iterations:
$\text{hp}_{\text{j}}=m+\text{HP-Score}_{\text{Dataset}}$ (7)
HP-Score for Adaptive HPF is:
@@ 3.4.3 HP-Score for Adaptive HPF
where $\text{hp}_{\text{j}}$ represents the HP-Score of sample $j$, x represents the level of HPF selected by prompt-selector at $i$th iteration at which task is addressed, $m$ represents the total number of levels in HPF, and $n$ denotes the total number of samples in the evaluation set.
4 Results
4.1 Results for Manual HPF
Table 2 presents the HP-Scores for the four LLMs: Llama 3 8B, Phi 3 3.8B, Mistral 7B, and Gemma 7B, across previously mentioned datasets with varying threshold values (0.15 and 0.20) for BLEU Papineni et al. (2002) and ROUGE Lin (2004). The BLEU and ROUGE thresholds at every prompting level serve as the task-solving criteria for each sample of IWSLT and SamSum respectively. The complexity at which the LLMs are able to handle the dataset effectively is reflected in the HP-Scores.
Figure 3: HP-Scores of all datasets and threshold values for LLMs in Manual HPF.
@@ 4.1 Results for Manual HPF
Llama 3 8B 1.3692 2.4281 3.3955 3.9246 1.3018 1.7166
Mistral 7B 1.4063 2.4929 3.5183 4.1455 1.2590 1.6770
Phi 3 3.8B 1.3965 2.1714 3.8479 4.4401 2.0925 3.4010
Gemma 7B 1.4558 2.5053 5.3929 5.8369 3.3137 5.0316
Table 2: HP-Scores of datasets (with thresholds) across all LLMs for Manual HPF.
Dataset Metric Threshold Llama 3 8B Phi 3 3.8B Mistral 7B Gemma 7B
BoolQ Accuracy - 0.99296 0.98685 0.98073 0.99419
CSQA Accuracy - 0.84766 0.88452 0.82063 0.83783
IWSLT BLEU 0.15 0.23539 0.19327 0.21629 0.08049
0.2 0.23091 0.19098 0.21569 0.06807
SamSum ROUGE-1 0.15 0.29395 0.21389 0.28529 0.16067
0.2 0.29996 0.22109 0.29365 0.14063
Table 3: Evaluation scores for datasets across all LLMs for Manual HPF.
Figure 3 shows that Llama 3 8B achieves lower HP-Scores than other LLMs across all datasets, indicating superior performance and the ability to solve tasks with simpler prompting strategies. All four LLMs were better at solving BoolQ than humans, with lower HP-Scores than BoolQ HP-Score by human experts. IWSLT was the task where all four LLMs were incapable of solving even at threshold values of 0.15 and 0.20, indicating that LLMs are incapable in solving domain-specific tasks compared to human experts in the respective domains.
Table 3 shows the evaluation scores for each LLM across all datasets using the Manual HPF. Llama 3 8B and Gemma 7B outperform Phi 3 3.8B and Mistral 7B on BoolQ, with all four LLMs achieving near-perfect accuracy. Phi 3 3.8B outperformed all other LLMs in solving CSQA. Despite having the fewest parameters, phi 3 3.8B achieved the lowest HP-Score among all four LLMs while also achieving the highest accuracy. Llama 3 8B has the highest BLEU scores and ROUGE scores across both threshold values for IWSLT and SamSum datasets respectively. While Gemma 7B has the lowest performance and the highest HP-Scores for the IWSLT and SamSum datasets. Llama 3 8B and Mistral 7B are the most capable and genearlized LLMs, achieving superior performance across all datasets. Phi 3 3.8B, even with the fewest parameters, was able to compete with Llama 3 8B and Mistral 7B. Gemma 7B even though was able to solve BoolQ and CSQA with superior performance showing it’s reasoning abilities but was unable to achieve good performance across domain-specific datasets like IWSLT and SamSum.
Model BoolQ CSQA IWSLT (0.15) IWSLT (0.20) SamSum (0.15) SamSum (0.20)
Llama3 5.2173 5.9136 6.2006 6.2841 5.0316 5.5756
@@
Table 4: HP-Scores of datasets (with thresholds) across all LLMs for Adaptive HPF.
Dataset Metric Threshold Llama3 8B Phi 3 3.8B Mistral 7B Gemma 7B
@@
Table 5: Evaluation scores for datasets across all LLMs for Adaptive HPF.
4.2 Results for Adaptive HPF
The Adaptive HPF automates the selection of appropriate prompting strategy for each task using a prompt-selector utilizing Llama 3 8B for the experiments. Table 4 shows the HP-Scores and Table 5 presents the evaluation scores for the LLMs using the adaptive HPF.
Figure 4: HP-Scores of all datasets and threshold values for LLMs in Adaptive HPF.
The findings presented in Figure 4 show that, on all datasets, the adaptive HPF produces HP-Scores for the LLMs that are higher than those obtained with the manual HPF, with decreased evaluation scores. Higher HP-Scores demonstrate that adaptive HPF is obviously unable to complete a variety of tasks, with the primary cause being hallucinations when choosing the appropriate level of prompting by the prompt-selector. Hallucination of prompt-selector leads to repeatedly choosing irrelevant prompting levels or choosing levels out of the range of manual HPF for several iterations of adaptive HPF, resulting in high HP-Scores. This demonstrates that LLMs are incapable of judging their evaluations of various tasks.
Figure 5: Accuracy scores of BoolQ and CSQA datasets across all LLMs for manual HPF and Adaptive HPF.
Figure 6: BLEU scores of IWSLT dataset across all LLMs for manual HPF and Adaptive HPF..
Figure 7: ROUGE-1 scores of SamSum dataset across all LLMs for manual HPF and Adaptive HPF..
4.3 Comparison between Manual HPF and Adaptive HPF
Figure 5 illustrates that the manual HPF outperforms the adaptive HPF on the BoolQ dataset across all LLMs. The decrease in accuracy with the adaptive HPF demonstrates its inability to select the most appropriate prompting level within its iteration limit. Figure 5 shows a significant decline in performance for LLMs in solving CSQA, suggesting that manual HPF is more dependable than adaptive HPF for solving CSQA. Despite solving the IWSLT with slightly worse performance as shown in Figure 6 when using the adaptive HPF, LLMs are unable to identify the optimal prompting level indicated by their high HP-Scores. Figure 7 shows that both frameworks perform similarly in solving SamSum, with the adaptive HPF showing higher HP-Scores across all LLMs. Table 1 shows that the HP-Scores of both manual HPF and adaptive HPF indicate that manual HPF more closely matches human estimation of task complexity, while the adaptive HPF faces difficulty in solving these tasks when compared to both manual HPF and humans. Both frameworks have higher HP-Scores when solving IWSLT than domain experts’ HP-Scores, indicating that pre-trained LLMs with more domain-specific knowledge are needed to demonstrate better performance across domain-specific tasks with less complex prompting strategies, reducing the need for these LLMs to be fine-tuned. Another important finding from the IWSLT and SamSum evaluation scores is that all four LLMs were unable to perform above a particular evaluation score at higher threshold values, as evidenced by the saturation of their performance across all threshold values.
Overall, our experiments demonstrate that different tasks don’t always require sophisticated prompting strategies from LLMs; it largely depends on the LLM’s aptitude for the task. The HP-Score provides a universal measure of task complexity relative to different agents, offering insights into the capabilities of both LLMs and humans. When comparing Manual HPF and Adaptive HPF, the Manual HPF is more reliable and consistently performs well on datasets with lower HP-Scores. In contrast, the Adaptive HPF is computationally inefficient due to hallucinations when selecting the prompting level in least number of iterations. For other tasks and LLMs, Adaptive HPF requires finetuning of the prompt selector or other techniques to minimize hallucinations and ensure consistent outcomes.
5 Conclusion
The HPT, combined with the HPFs, offers a robust and efficient method for evaluating LLMs based entirely on task complexity. These findings highlight the effectiveness of the HPF in evaluating the task with the most effective prompting strategies based on task complexity, leading to enhanced LLM performance across various datasets and tasks. This approach not only provides a detailed understanding of LLMs problem-solving capabilities but also introduces Adaptive HPF by dynamically selecting the most appropriate prompting strategy, paving the way for future advancements in LLM evaluation techniques.
6 Limitations
Restricted Model Evaluation: The study focused on four specific instruction-tuned LLMs: Llama 3 8B, Phi 3 3.8B, Mistral 7B, and Gemma 7B. While these LLMs provide valuable insights, the inclusion of a more diverse set of LLMs, both in terms of architecture and scale, would help assess the applicability of the HPT across a broader spectrum of LLMs.
Restricted Dataset Evaluation: The evaluation of both HPF and adaptive HPF is conducted on four distinct datasets encompassing tasks such as question-answering, reading comprehension, translation, and summarization. Incorporating a wider variety of datasets that cover additional tasks would aid in evaluating the HPT’s applicability across a more extensive range of tasks.
Manual framework optimization: The effectiveness of the HPF heavily relies on the quality of the prompts used at each level of the taxonomy. Crafting high-quality prompts that accurately reflect the subtleties of each level demands considerable expertise and repeated refinement. This study only investigated a limited set of prompting strategies within the HPF, indicating a need for further research into creating diverse structural frameworks and incorporating additional prompting strategies.
Adaptive framework optimization: The Adaptive HPF automates the selection of suitable prompting strategies according to task complexity. However, the current system depends on a prompt-selector that utilizes an LLM. There were numerous hallucinations when choosing the appropriate prompting strategy for the task. Further investigation is required to minimize hallucinations in the prompt-selector and to explore alternative methods, such as ensemble techniques or meta-learning approaches, to enhance the efficiency and effectiveness of the adaptive framework.
7 Ethical Consideration
The HP-Scores assigned by experts in the domains related to the datasets—BoolQ, CSQA, IWSLT and SamSum—may introduce bias into the comparative analysis. This potential bias stems from the subjective nature of expert scoring, which can be influenced by individual experience and perspective. Despite this, the datasets utilized in this study are publicly available and widely recognized in the research community, thereby minimizing the risk of unanticipated ethical issues. Nevertheless, it is crucial to acknowledge the possibility of scoring bias to ensure transparency and integrity in the analysis.
Appendix A Large Language Models Used for Evaluation
@@
The HPF supports multiple LLMs and includes mechanisms for optimizing performance through advanced quantization techniques. The experiments were conducted on the following instruction-tuned LLMs and the model description and licenses are discussed in Table 6.
@@
Llama 3 8B Research License Non-commercial, requires attribution, ethical guidelines
Phi 3 3.8B Open-source Broad use allowed, must include original license and notices
@@
Gemma 7B Open-source/Research Depending on license, may have non-commercial restrictions or broad use allowed
@@
The selection of Llama 3 8B, Phi 3 3.8B, Mistral 7B, and Gemma 7B offers a comprehensive evaluation set that spans different model sizes, architectures, and use cases, ensuring the HPF framework is robust and generalizable. Llama 3 8B excels in language understanding and generative tasks, achieving top-tier performance across benchmarks such as MMLU, HumanEval, GSM-8K, BoolQ, and CommonSenseQA. This model’s proficiency in diverse tasks, including complex mathematical reasoning and commonsense reasoning, underscores its high-quality text generation capabilities. Phi 3 3.8B, optimized for environments with limited computational resources, performs competitively on benchmarks like GSM-8K and BoolQ, making it ideal for real-time applications due to its efficiency and quick response times. Mistral 7B balances parameter count and performance, excelling in MMLU, CommonSenseQA, and HumanEval, showcasing versatility in language understanding and code-related tasks. Gemma 7B, specialized for particular domains, excels in niche areas like medical text understanding and technical documentation. While not always the top performer in general benchmarks, Gemma 7B’s specialization complements the general-purpose nature of the other LLMs, ensuring the HPF framework is applicable to a wide range of scenarios and tasks.
By including these LLMs, the effectiveness of the HPF is highlighted, demonstrating its ability to accommodate a wide range of scenarios and tasks, from general-purpose language understanding and generation to specialized domain-specific applications. This comprehensive evaluation ensures that HPF is a robust and generalizable framework in the landscape of LLM evaluation and prompting frameworks.
The quantization configuration for each of the instruction-tuned LLM involves reducing the precision of the LLM’s weights and computations, allowing for a more compact representation and faster processing. The LLMs were loaded in a 4-bit precision format which significantly reduces the memory footprint without compromising accuracy. By using half-precision floating point computations (float16), the LLMs maintain high computational efficiency. Additionally, the quantization type employed is NF4, which provides a balanced approach between model size and performance. The use of double quantization further refines this process, ensuring that the LLMs operate efficiently even on resource-constrained hardware.
The generation configuration is carefully designed to optimize the quality of text output during evaluation. This involves setting parameters that influence the behavior of the LLMs during the text generation process. For instance, the maximum number of tokens that can be generated in a single run is capped (1024 tokens) to ensure concise and relevant outputs. The temperature parameter was set to 0.6 to control the randomness of the predictions; lower temperature results in more deterministic and focused outputs. Top-p (p=0.9) sampling allows the LLMs to explore various possible continuations for a given input. Finally, the repetition penalty is applied to discourage the LLMs from repeating the same phrases or sentences, thus ensuring more coherent and varied text generation.
Appendix B Prompt Templates
B.1 Level 1: Role Prompting
Role prompting is the most basic level of interaction with a LLM, where the LLM is given a specific role or task to perform without additional context or examples. This technique relies solely on the initial instruction to guide the LLM’s response. For instance, if the LLM is asked to "act as a translator", it will attempt to translate text based on its training data. This method is straightforward but may lack depth and context, leading to less accurate or nuanced outputs. Table 7 presents the prompt templates utilized for Role prompting across all datasets in our experiments.
@@ B.1 Level 1: Role Prompting
B.2 Level 2: Zero-shot Chain-of-Thought Prompting
Zero-shot Chain-of-Thought prompting builds upon basic role prompting by asking the LLM to generate a reasoning process for a given task, even though it has not been explicitly trained on similar examples. In this approach, the LLM attempts to break down the problem and solve it step-by-step using its internal knowledge base. This technique leverages the LLM’s ability to infer and reason, improving the quality of the response by encouraging logical progression and coherence in the thought process. Table 8 presents the prompt templates utilized for Zero-CoT across all datasets in our experiments.
@@ B.2 Level 2: Zero-shot Chain-of-Thought Prompting
B.3 Level 3: Three-Shot Chain-of-Thought Prompting
Three-shot Chain-of-Thought prompting advances the zero-shot approach by providing the LLM with three examples of the task at hand, including the reasoning steps taken to arrive at the solution. These examples serve as a guide, helping the LLM understand the structure and logic required to solve similar problems. By seeing how previous tasks were approached and resolved, the LLM can better mimic the process, leading to more accurate and contextually appropriate responses. Table 9 presents the prompt templates utilized for 3-CoT across all datasets in our experiments.
@@ B.3 Level 3: Three-Shot Chain-of-Thought Prompting
Table 9: Prompt templates of different datasets for Three-Shot Chain-of-Thought Prompting
B.4 Level 4: Least-to-Most Prompting
Least-to-most prompting is a sophisticated technique that involves gradually increasing the complexity of prompts presented to the LLM. The process begins with simpler, more manageable tasks and progressively moves to more complex challenges. This method helps the LLM build confidence and competence in solving problems, as it can leverage the insights gained from simpler prompts to tackle more difficult ones. It enhances the LLM’s ability to generalize from straightforward examples to more intricate scenarios. Table 10 presents the prompt templates utilized for Least-to-Most Prompting across all datasets in our experiments.

### v4

@@ E.1 HPI for Adaptive HPF
for sample $j$ in the evaluation dataset do
@@
     for iteration $i=1$ to $m$ do
         Select prompting strategy at level $x$
         if LLM completes the task at iteration $i$ then
[…]
              break
         end if
     end for
     if solved = False then
[…]
     end if
end for
$\texttt{HPI}_{\texttt{Adaptive}}=\frac{1}{n}\sum_{j=1}^{n}\texttt{HPI\_List}[j]$
$x$ represents the level of the HPF selected by prompt-selector at $i$th iteration at which the task is addressed, $m$ represents the total number of levels in the HPF, and $n$ denotes the total number of samples in the evaluation set.
@@
Llama-3 8B 5.2173 5.9136 6.2006 6.2841 5.0316 5.5756
@@
Table 6: HPI (lower is better) of LLMs across datasets (with thresholds) for Adaptive HPF.
Dataset Metric Threshold Llama-3 8B Phi-3 3.8B Mistral 7B Gemma 7B
@@
Table 7: Performance scores of LLMs across datasetsfor Adaptive HPF.
E.2 Hallucination in Adaptive HPF
Hallucinations in prompt-selector refer to instances where the LLM generates incorrect or misleading prompting levels or nonsensical information that is not supported by the HPF. These hallucinations can occur across various tasks, including question answering, multiple-choice questions, translation, and summarization.
For the BoolQ task, the prompt-selector initially struggles, indicated by the iterations where it reaches Level 4 with hallucinations (represented by ’…’). However, by the fourth iteration, Llama-3 8B manages to answer correctly at Level 2. For the CSQA task, prompt-selector exhibits hallucinations initially, shown by Level 4 and Level 0 (not included in HPF) responses. Eventually, it corrects itself by the third iteration, providing the correct answer at Level 2. For the IWSLT task, prompt-selector demonstrates a consistent pattern of hallucinations across multiple iterations. Even though Llama-3 8B attempts the translation at Level 2 multiple times, it ultimately fails to provide a correct translation, indicating a persistent hallucination. For the SamSum task, prompt-selector shows initial hallucinations in the first three iterations (Level 4). However, by the fourth and fifth iterations, the prompt-selector starts producing lower levels. Finally, Llama-3 8B achieves the correct answer at Level 2 in the last iteration .
The results in Table 6 and Table 7 indicate that the prompt-selector exhibits hallucinations in selecting complexity levels across various tasks and iterations resulting in higher HPI for Adaptive HPF, with performance varying significantly. While the LLM can eventually produce correct answers, as seen in the BoolQ and SamSum tasks, it often requires multiple attempts and may still fail in tasks like IWSLT translation.
E.3 Prompt Template for Prompt-Selector
The prompt-selector in adaptive HPF selects the prompting level based on the task complexity to address the task. Llama-3 8B serves as the prompt-selector in the experiments. The prompt template was meticulously designed to ensure maximum clarity, aiming to reduce hallucinations and select the most effective prompting strategy.
  • Role Prompting: Defines a role for the model in solving the task.
  • Least-to-most prompting: Uses a sequential method to derive essential insights from the task to solve it.
Appendix F Computational Budget
All evaluation experiments and ablation studies were conducted on V100 GPUs (16GB and 32GB variants), utilizing a total of around 9,000 computation hours, this equates to approximately 1.125 petaflop-hours of computational resources.
Appendix G Large Language Models Used for Evaluation
@@
The HPF supports leading open source and proprietary LLMs and includes mechanisms for optimizing performance through advanced quantization techniques. The experiments were conducted on the following instruction-tuned LLMs, and the model description and licenses are discussed in Table 8.
@@
GPT-4o Proprietary Commercial use requires paid API access, subject to OpenAI’s terms of service
Claude 3.5 Sonnet Proprietary Commercial use requires paid API access, subject to Anthropic’s terms of service
Mistral-Nemo 12B Proprietary Usage likely restricted to authorized partners or specific use cases
Gemma-2 9B Research License Non-commercial use only, research purposes
Llama-3 8B Research License Specific restrictions may apply, typically for non-commercial research use
@@
Gemma 7B Open-source/Research Depending on the license, may have non-commercial restrictions or broad use allowed
Phi-3 3.8B Open-source Broad use allowed, must include original license and notices
@@
The LLMs were loaded in 4-bit precision format, with a maximum generation limit of 1024 tokens per run to ensure concise outputs. The temperature was set to 0.6 to control prediction randomness, while top-p sampling (p=0.9) enabled the exploration of diverse continuations. Additionally, a repetition penalty was applied to discourage the generation of repeated phrases, promoting coherent and varied text output.
Appendix H Prompt Templates
H.1 Level 1: Role Prompting
Role prompting represents the most basic interaction with an LLM, assigning it a specific role or task without additional context or examples. This approach relies solely on the initial instruction to guide responses. For instance, asking the LLM to “act as a translator” prompts it to translate text based on its training data. While straightforward, this method may lack depth, resulting in less accurate or nuanced outputs. Table 9 shows the prompt templates used for role prompting across all datasets in the experiments.
@@ H.1 Level 1: Role Prompting
GSM8k Based on the question: "question", calculate the numerical answer to the question as an expert mathematician.
HumanEval Complete the given code based on the mentioned constraints: "code" as an expert programmer.
MMLU Choose the answer: ‘‘question",A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4" as a critical thinker.
@@
H.2 Level 2: Zero-shot Chain-of-Thought Prompting
Zero-shot Chain-of-Thought (CoT) prompting enhances basic role prompting by requiring the LLM to generate a reasoning process for a task, despite not being explicitly trained on similar examples. This method encourages the LLM to break down the problem and solve it step-by-step using its internal knowledge, improving response quality through logical progression and coherence. Table 10 displays the prompt templates used for Zero-CoT across all datasets in the experiments.
@@ H.2 Level 2: Zero-shot Chain-of-Thought Prompting
GSM8k Based on the question: "question", calculate the numerical answer to the question. Let’s think step by step.
HumanEval Complete the given code based on the mentioned constraints: "code". Let’s think step by step.
MMLU Choose the answer: ‘‘question",A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4". Let’s think step by step.
@@
H.3 Level 3: Three-Shot Chain-of-Thought Prompting
Three-shot Chain-of-Thought (CoT) prompting builds on the zero-shot approach by providing the LLM with three task examples, including the reasoning steps used to reach the solution. These examples help the LLM grasp the required structure and logic, enabling it to better replicate the problem-solving process and produce more accurate, contextually relevant responses. Table 11 shows the prompt templates used for 3-CoT across all datasets in the experiments.
@@ H.3 Level 3: Three-Shot Chain-of-Thought Prompting
GSM8k Based on the question:"gsm8k_question1", calculate the numerical answer to the question.Answer: "gsm8k_ans1".Based on the question:"gsm8k_question2", calculate the numerical answer to the question.Answer: "gsm8k_ans2". Based on the question:"gsm8k_question3", calculate the numerical answer to the question.Answer: "gsm8k_ans3". Based on the question: "question", calculate the numerical answer to the question.
HumanEval Complete the given code based on the mentioned constraints: "humaneval_code1", Code: "humaneval_sol1". Complete the given code based on the mentioned constraints: "humaneval_code2", Code: "humaneval_sol1".Complete the given "code" based on the mentioned constraints: "humaneval_code3", Code: "humaneval_sol3".
MMLU Choose the answer for the question: "mmlu_ques1" A. [AND, NOT] B. [NOT, OR] C. [AND, OR] D. [NAND] Answer: C. Explanation: "mmlu_exp1". Choose the answer for the question "mmlu_ques2" A. The defendant’s statement was involuntary. B. The defendant’s statement was voluntary. C. The defendant was not in custody when the statement was made. D. The statement was not made in response to a known police interrogation. Answer: A, Explanation: "mmlu_exp2". Choose the answer for the question: "mmlu_ques3" . A. Wrong, Wrong. B. Wrong, Not wrong C. Not wrong, Wrong D. Not wrong, Not wrong. Answer: B Explanation: "mmlu_exp3". Choose the answer."question" ‘‘question",A. ‘‘option 1",B. ‘‘option 2",C. ‘‘option 3",D. ‘‘option 4".
Table 11: Prompt templates of different datasets for Three-Shot Chain-of-Thought Prompting.
H.4 Level 4: Least-to-Most Prompting
Least-to-most prompting is an advanced technique that gradually increases prompt complexity, starting with simpler tasks and progressing to more complex challenges. This method allows the LLM to build confidence and leverage insights from easier prompts to tackle harder ones, enhancing its ability to generalize from straightforward examples to intricate scenarios. Table 12 displays the prompt templates used for Least-to-Most Prompting across all datasets in the experiments.